OpenStudy (anonymous):

Find the fourth roots of the complex number " z1 = 1 + √3 * i

OpenStudy (anonymous):

Part I: Write z1 in polar form. Part II: Find the modulus of the roots of z1. Part III: Find the four angles that define the fourth roots of the number z1. Part IV: What are the fourth roots of the equation " z1 = 1 + √3 * i ".

OpenStudy (anonymous):

Is my answer correct? a = 1 b = sqrt(3) sqrt(a^2 + b^2) = sqrt(1 + 3) = sqrt(4) = 2 z = 2 * (1/2 + i * sqrt(3)/2) z = 2 * (cos(pi/3 + 2pi * k) + i * sin(pi/3 + 2pi * k)) z = 2 * (cos((pi/3) * (1 + 6k)) + i * sin((pi/3) * (1 + 6k))) z^(1/4) = 2^(1/4) * (cos((pi/12) * (1 + 6k)) + i * sin((pi/12) * (1 + 6k))) 2^(1/4) * (cos(pi/12) + i * sin(pi/12)) 2^(1/4) * (cos(7pi/12) + i * sin(7pi/12)) 2^(1/4) * (cos(13pi/12) + i * sin(13pi/12)) 2^(1/4) * (cos(19pi/12) + i * sin(19pi/12))

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

modulus is 2

OpenStudy (anonymous):

angle is \(\frac{\pi}{3}\)

ganeshie8 (ganeshie8):

you may verify the answers by rising the roots to 4th power you should get back the z1

OpenStudy (anonymous):

and \[\frac{\pi}{3}\times \frac{1}{4}=\frac{\pi}{12}\]

OpenStudy (anonymous):

all looks swell

OpenStudy (anonymous):

Should I simplify the roots?

OpenStudy (anonymous):

@satellite73

ganeshie8 (ganeshie8):

i don't think you can simplify them further

OpenStudy (anonymous):

I meant the angles.

ganeshie8 (ganeshie8):

they look good the way they are now

ganeshie8 (ganeshie8):

pi/12 what can you simplify here ?

OpenStudy (anonymous):

@satellite73 said it simplifies to pi/3?

ganeshie8 (ganeshie8):

Nope, satellite was referring to something else

OpenStudy (anonymous):

ok